Help me with the workings for questions 3 and 4 please !!

Question 3 Not answered Marked out of 1.00 P Flag question A linear programming problem seeks to maximise R where R- 2.8x+y subject to x y 300 x 2.6 xy 20 What is the maximum value of R? Give your answer correct to two decimal places.) Answer: The correct answer is: 304.68 Question 4 Not answered Marked out of 1.00 A couple are planning for their retirement. The chance that Arthur will survive for ten years is 0.82. The chance that Jenny will survive for ten years is 0.71. What is the chance that exactly one of them will survive for ten years? Give you answer correct to four decimal places.) Flag question ( Answer: The correct answer is: 0.37
Answer #1

3)

It has to be solved using simplex method or graphical method. You will need to learn Linear Programming for that.

Having said that, the problem in hand is simple and we can solve it through logical thinking.

R = 2.8x + y

So, we can understand that for any increase in x, there will be 2.8 times increase in R whereas, for an increase in y, there will be a similar increase in R.

ty300

So, every increase in x, y has to be decreased in same quantity and vice versa. Thus, if we increase x by one unit, we decrease y by one unit and the net increase in R will be 2.8 - 1 = 1.8 units. So, to maximise R, we need to maximise x.

But, x is constrained by,

xleq 2.6

That is, 2.6 is the maximum value x can take. Therefore,

У 300-2.6 = 297.4

R = 2.6*2.8 + 297.4 = 304.68

4)

We need to find the probability of exactly one person surviving.

extup{P(exactly one surviving) = P(only Arthur survives) + P(only Jenny survives) }

extup{P(only Arthur survives)= P(Arthur survive) * P (Jenny do not survive)}

extup{P(only Arthur survives)}= 0.82 * (1-0.71) = 0.82 * 0.29 =0.2378

extup{P(only Jenny survives)= P(Jenny survive) * P (Arthur do not survive)}

extup{P(only Jenny survives)}= 0.71 * (1-0.82) = 0.71 * 0.18 = 0.1278

herefore extup{P(exactly one surviving)} = 0.2378 + 0.1278 = 0.3656 approx 0.37